OpenStudy (anonymous):

Water freezes at thirty-two degrees Fahrenheit, or zero degrees Celsius. Use temperature conversion formulas to show that these two measures are indeed the same temperature.

OpenStudy (anonymous):

The conversion formula looks like this: Tc = (5/9)*(Tf-32); where Tc = temperature in celcius and Tf = Temperature in Fahrenheit. so if we fill it in it looks like this: (0) = (5/9)*(32-32) Then we work it out. (0) = (5/9)*0 0 = 0 and there's your answer 32 degrees Fahrenheit equals zero degrees Celsius.

OpenStudy (anonymous):

this is a question to were u have to show ur work....